What is the probability that an electron will tunnel through a 0.54 nm gap from a metal to a STM probe if the work function is 4.2 eV ?
Answer #1

Electron's penetration distance is given by:

d = h(bar)/sqrt (2*m*(U0 - E))

U0 - E = Work function = 4.2 eV

So,

d = 1.05*10^-34/sqrt (2*9.11*10^-31*4.2*1.6*10^-19)

d = 9.49*10^-11 m = 0.0949 nm

Now tunnelling probability through barrier is given by:

Ptunnel = exp(-2w/d)

Ptunnel = exp (-2*0.54/0.0949) = 1.142*10^-5